# Integrating Multi-Physics Simulations and Machine Learning to Define the Spatter Mechanism and Process Window in Laser Powder Bed Fusion

### v1

@@ 3.2 Spatter and Meltpool Property Correlations
Meltpool dynamics relies on the process parameter, and spatter ejected from the meltpool can have properties close to the meltpool; however, it is unknown how close this properties can be due to unavailability of equipment that has the tolerance and sensitivity to detect the variations. Collecting measurements during the vapor plume generation from the meltpool further complicates the data extraction due to the poor visibility and short timescale between changing meltpool dynamics and spatter ejections. Harnessing the strength of high fidelity 3D simulation and ML models, the slight variation in the properties can appear in patterns across measurements and be detected; aiding in differentiating the measured properties of both the meltpool and the spatter at ejections. Informative statistical distribution was generated as shown in the Figure 4 and Figure 5. The diagonal of the pairplot displays a plotted histogram showing the distribution of the property measurements. On the left side of the diagonal, are the Kernel Density Estimation (KDE) plots showing the data density concentration while on the right-hand-side of the diagonal histogram are the scatter plots of the distribution. The histogram plots along the diagonal suggest that spatter has a wider spatial distribution, with its density spreading out over a larger range. Meltpool, on the other hand, shows a more concentrated spatial distribution, indicating a more localized phenomenon. Additionally, the wider distribution with respect to the spatial distributions (x, y, z) can be attributed to the possibility of the spatter ejection being detected within and away from the meltpool location as shown in the Figure 4 which can depend on ejection velocity and simulation’s timestep. The wider spatial distribution of the spatter is more prominent along the y-axis. The circular shape of the laser beam induces curvature growth of the meltpool, such that its breakage will lead to spatter detection in larger different between the magnitude of its standard deviation and the meltpool than what was observed for other axes. A wider separation was observed between the histogram plot of the meltpool and the spatter along the z-axis, and it can be attributed to the spatter ejection mostly occurring and detected at a higher height than the meltpool. For the velocity distributions as shown on Figure 5, except for y-velocity, the Spatter KDE plots, histogram and scatter plots are broader, suggesting a wider range of velocities, while the Meltpool plots are narrower, indicating more uniform velocities. y-velocity distributions appear to be similar with respect to the spatter and the meltpool, suggesting that y-velocity might be a weak parameter in driving spatter ejection. Also, density distributions between the spatter and the meltpool appear to be close to each other, with the spatter possessing slightly broader distribution to the upper end, indicating evidence of cooling. In general, spatter data points on the KDE plots, histogram and the scatter plots exhibit a wider spread, reinforcing the idea that spatter exhibits a greater variability in spatial and velocity distributions, while meltpool shows tighter clustering.
@@
In contrast, wider distribution spreads were observed for the temperature and pressure of the meltpool in comparison to the spatter, implying a broader temperature and pressure range of the meltpool. Spatter shows higher peaks in the KDE plots, indicating temperatures are more consistent and concentrated around a mean value. The scatter plots for the spatter display more defined groupings, while the meltpool’s points are dispersed over a wider range, confirming the trends suggested by the KDE plots. Meltpool’s pressure distribution is wider also as indicated by its KDE plot, whereas spatter exhibits a sharper peak, suggesting a narrower range of pressure values is required for its ejections. The scatter plots corresponding to the pressure of the meltpool against other parameters show a more dispersed pattern. The wider temperature and pressure distribution of the meltpool dataset is due to the high gradients of these properties along the meltpool volume at a given timestep. Simultaneously, the narrower distribution of the spatter’s temperature and pressure in addition to it less concentration at the upper bound in comparison to the meltpool could be due to it ejection occurring away from the laser impact on the material and occurrence of cooling, which could depend on the offshoot velocity that will determine how far away it is from the laser. Overall, spatter represents a more variable and less controlled process with a broader range of values across most of the parameters, whereas meltpool indicates a more stable and consistent process where the parameters are more narrowly distributed around certain central values. Spatter ejection from the meltpool could be attributed to the locations in the meltpool that possess the observed magnitude or threshold for the combination of parameters. However, it is crucial to quantify these parameters based on their importance to spatter generation.
@@
Figure 8: Spatter prediction on FLOW-3D simulation generated from testing FLOW-3D dataset on ML model that was trained using OpenFOAM. Region masked with black are identified as location predicted as spatter (a) prediction at timestep 65 $\mu s$ (b) prediction at timestep 165 $\mu s$ (c) prediction at timestep 265 $\mu s$.
3.3 Spatter and Meltpool Classification via ML Model
Classification task was conducted using the following ML models; Extremely Randomized Trees(ExtraTrees), Extreme Gradient Boosting(XGB), LGBM(Light Gradient Boosting Machine), RF(Random Forest), Bagging, Support Vector Classifier(SVC) and k-nearest neighbors(KNN). Our dataset was extracted from OpenFOAM simulation of two experiments at 545 W with scan speed of 2 m/s and 690 W and scan speed of 2 m/s. The dataset consists of 358 points whereby 50 % is spatter and the remaining 50 % is associated with the meltpool. The train-test split used is 70 %-30 %. Upon screening different ML models, where the collected features include x-axis, y-axis, z-axis, temperature, velocity, x-velocity, y-velocity, z-velocity, density and pressure, high accuracy were observed for all the ML models, with ExtraTrees being the highest at 96 % and KNN being the lowest at 94 % as shown on the figure 7a. The importance of the features used in ML testing, as shown on the Figure 7b, followed the trend: z-axis $>$ pressure $>$ z-velocity $>$ y-axis $>$ temperature $>$ velocity $>$ x-velocity $>$ y-velocity $>$ density. The highest contribution of the z-axis to the accuracy is attributed to the wide separation observed between the spatter and the meltpool when z-axis is paired together with itself, or other parameters as shown on the histogram, KDE, and scatter plots as shown on Figure 4.
@@ 3.3 Spatter and Meltpool Classification via ML Model
The methodologies of our data extraction and combination of parameters for this classification task is promising. These techniques can serve as a benchmark in predicting spatter or developing reduced order model package for spatter generation either with desired simulation packages or collected experimental measurements. Consequently, the high accuracy outcome of ML predictions underscore the efficacy of our ML models in predicting spatter events on the commercial software package, FLOW-3D, which lacks the capability to accurately simulate spatter. Therefore, modelling LPBF in FLOW-3D generates a dataset that is specifically for the meltpool with no spatter occurrence. Although some locations in the meltpool dataset generated by the FLOW-3D could be having spatter-like properties or the property-threshold required for spatter ejection, but rather remain in the meltpool due to the absence of physics module required to resolve the mass transfer between the interface of the different phases. Therefore, adapting the current ML model that was trained on OpenFOAM dataset will necessitate removal of spatial components (x, y, z) from the features that will be used to train and test ML models. On training and testing OpenFOAM dataset without spatial components, high accuracies were still achieved with slight decrease. Both ExtraTrees and RF exhibit highest values at 93 % while KNN has the lowest value of 86 % as shown in Figure 7b. It was noted that RF changed the least by 0.9 %, while ExtraTrees changed by 1.7 % and KNN changed the most by 8 % as compared to their respective initial accuracy that were achieved with spatial component being included in the features. Consequently, RF was chosen for predicting spatter on FLOW-3D software due to its best stability as it demonstrated least changes with different feature configurations. The trend observed in the feature importance studies without spatial components, as shown in Figure 7d follows the trend of pressure $>$ z-velocity $>$ temperature $>$ velocity $>$ x-velocity $>$ y-velocity $>$ density. Evidently, pressure is not only the most significant parameter driving the dynamics in the meltpool, but also dictates possibility of spatter ejection as its magnitude is confined within a narrow range as shown in Figure 5. Strong importance observed with respect to z-velocity can be understood from the tendency of a body that possesses high z-velocity to travel upward, and we observed a narrow range for meltpool z-velocity, which is a subset of spatter’s z-velocity. Also, observing temperature having less importance than the z-velocity despite its direct relationship with pressure according to Clausius–Clapeyron relation could be attributed to the saturation that is reached such that increment in the temperature will result in boiling of the material which will not lead to spatter ejection, but more plume generation. High contribution of z-velocity to the spatter ejection could also be understood from the parabolic shape of the meltpool which is formed as the laser traverse in x-axis direction. Hot fluid will build up in the direction of the laser, which can break up to create spatter due to its elongation as shown in figure 6a, via shear thinning. Additionally, spatter breakage can further intensify the backward dispersion of material and energy that are in form of a wave which can travel through the high gradient region of the meltpool, creating z-velocity component that could enhance ejection. The backward dispersion of material and energy in the meltpool can be attributed to the combination of other factors such as pressure, temperature gradient, and the vibration in the meltpool which can also influence other properties like viscosity and surface tension to cause Marangoni effect and promote meltpool flow, elongation and ultimately fluid ejections.
@@
To gain further insight to the distribution of ML model predictions, we constructed a precision recall table as shown on Figure 7e and 7f for predictions with spatial components included in the features and with no spatial components respectively. We observed our model having the capability in predicting most of the points with the meltpool and the spatter. The most wrong prediction observation for the model with the presence of spatial component features and non-spatial component features is associated with regions that were meltpool but rather predicted as spatter. Upon testing our model on FLOW-3D, voxels that matches the property of spatter were identified and masked with black points as shown on Figure 8. At the early timestep of the FLOW-3D simulation, spatter predictions were predominantly observed at the front of the meltpool (Figure 8a), then at the latter timesteps (Figure 8a and the 8b), spatter tends to occupy the backward side of the meltpool, this is similar to observed OpenFOAM simulation for spatter generation as shown in Figure 6. It should be noted that conditions for spatter generation were not met at the location of the hotspot of the meltpool due to the high temperature that could be causing fluid evaporation. Consequently, the high accuracy outcome of RF underscores the efficacy of our developed methodology for the reduced-order task on spatter prediction whereby testing datasets of different process parameters generated from commercial software, FLOW-3D towards development of process map is possible.
@@ 3.4 Spatter Process Map
Figure 9: Process map generated via hybrid of OpenFOAM trained ML model on FLOW-3D dataset to understand relationship between spatter formation and process conditions. The process map was created using 283 FLOW-3D simulations with different combinations of power and velocity
Display of spatter quantification with varying power/velocity combination will provide access to suitable operating window with minimal defects for producing part. Running openFOAM for additive manufacturing is computationally expensive due to many physics it is configured to solve as compared to FLOW-3D, a devoted package for AM, but lacks physics to solve spatter generation. On the same machine, FLOW-3D is 100 times faster than OpenFOAM in generating LPBF solution. Building hybrid of OpenFOAM and FLOW-3D via the high accuracy ML model as a reduced order model is attractive for large dataset generation for the development of spatter process map. As shown in Figure 9, the spatter quantification follows a predictable trend within processing parameter space for laser LPBF. It was observed that more spatter was generated at high power low velocity which is the region observed for keyhole formation in porosity. At low power and high speed, less spatter was observed and this can be associated with the region know as lack of fusion [12]. The reason for the less spatter at this region can be attributed to less interaction of the material and the laser, creating less melting resulting in the melt track that will not be fully join. Based on the operating conditions, we observed the trend follow by the quantity of spatter generation as high power/low speed $>$ high power/high speed $>$ low power/low speed $>$ low power/high speed. Safe operating window lies between the region of lack of fusion and keyhole. Although the safe operating might possess slightly more spatter than the lack of fusion region, however as it has been previous studied than operating at this region prevents defects associated with porosity, additionally machine improved design or shielding gas flow could aid spatter elimination while operating within the safe operating window.

### v2

@@ 3.2 Spatter and Melt Pool Property Correlations
The behavior of the melt pool depends on the process parameters, and while the ejected spatter originates from the melt pool, likely its properties are different. However, the exact nature and extent of these differences remain unknown as a result of the lack of equipment with the precision and sensitivity to detect such subtle variations. Collecting measurements during the vapor plume generation from the melt pool further complicates the data extraction because of the poor visibility and short timescale between changing melt pool dynamics and spatter ejections. Harnessing the strength of high-fidelity 3D simulation and ML models, the slight variation in the properties can appear in patterns across measurements and be detected; aiding in differentiating the measured properties of both the melt pool and the spatter at ejections. Statistical distributions were generated to demonstrate this effect, as shown in Figures 7 and 8. The diagonal of the pairplot displays a plotted histogram showing the distribution of the property measurements. On the left-hand side of the diagonal are the kernel density estimate (KDE) plots showing the data density concentration, while on the right-hand side of the diagonal histogram are the scatter plots of the distribution. The histogram plots along the diagonal for most of the measured parameters suggest that the spatter has a wider spatial distribution, with its density spread over a larger range. The melt pool, on the other hand, shows a more concentrated spatial distribution, indicating a more localized phenomenon. A wider distribution was observed with respect to the spatial distributions (x, y, z) of the spatter compared to that of the melt pool on the histograms along the diagonal of the plot in Figure 7. The wider distribution can be attributed to the possibility that the ejection of the spatter is detected within and away from the location of the melt pool, which can depend on the velocity of the ejection and the simulation timestep. The wider spatial distribution of the spatter is more prominent along the y-axis. The circular shape of the laser beam can induce curvature growth of the melt pool, which, upon breaking, leads to spatter ejection and detection in the gas phase. The distance between the melt pool and the most ejected spatter is relatively short when compared to the size of the x-axis (laser traveling direction) and z-axis (simulation height) as shown in Figure 2. The scatter plot supports the prominence along the y-axis compared to the x and z axes. Consequently, a larger portion of the y-axis domain is likely to see the presence of spatter compared to the x-axis and the z-axis. A wider separation was observed between the histogram plot of the melt pool and the spatter along the z-axis, and it can be attributed to the spatter ejection being mostly occurring and detected at a height higher than that of the melt pool. For the velocity distributions shown in Figure 9, except for the y velocity, the KDE spatter plots, the histogram, and the scatter plots are broader, suggesting a wider range of velocities, while the melt pool plots are narrower, indicating more uniform velocities. y-velocity distributions appear to be similar with respect to the spatter and the melt pool, suggesting that y-velocity might be a weak parameter in driving spatter ejection. In addition, the density distributions between the spatter and the melt pool appear to be close to each other, with the spatter possessing a slightly broader distribution to the upper end, indicating evidence of cooling. Small amounts of spatter fall into this category, as shown in the scatter plot of density with respect to other parameters, indicating that most of the spatter are in the molten state as the melt pool. In general, spatter data points on the KDE plots, histogram, and scatter plots exhibit a wider spread, reinforcing the idea that spatter exhibits greater variability in spatial and velocity distributions, while melt pool shows tighter clustering.
@@
In contrast, wider distribution spreads for the temperature and pressure of the melt pool were observed compared to those observed for the spatter, implying a broader temperature and pressure range of the melt pool. Spatter shows higher peaks in the histogram plots, indicating that the temperatures are more consistent and concentrated around a mean value. The scatter plots for the spatter display more defined groupings, while the melt pool’s points are dispersed over a wider range, confirming the trends suggested by the histogram plots. The melt pool pressure distribution is wider also, as indicated by its histogram and scatter plot, whereas the spatter exhibits a sharper peak. This suggests a narrow range of pressure values could be required for spatter ejections, assuming that the melt pool’s pressure is close to the spatter’s pressure at the point of ejection given the evidence of pressure overlap on the histogram plot. The scatter plots corresponding to the pressure of the melt pool against other parameters show a more dispersed pattern. The wider temperature and pressure distribution of the melt-pool data set is due to the high gradients of these properties along the melt-pool volume. Simultaneously, the narrower distribution of the spatter’s temperature and pressure along with the reduced concentration at the upper bound compared to the melt pool could be due to its ejection occurring away from the laser impact on the material, resulting in cooling. The magnitude of cooling depends on the velocity of the ejection, especially when it travels away from the laser. In addition, the ejected spatter experiences pressure forces in all directions in the gas phase, causing an increase in the surface tension and the sphericity of their shape. Overall, the spatter represents a more variable and less controlled process with a broader range of values across most of the parameters, whereas the melt pool indicates a more stable and consistent process where the parameters are more narrowly distributed around certain central values. Spatter ejection from the melt pool could be attributed to the locations in the melt pool that possess the observed magnitude or threshold for the combination of parameters. However, it is crucial to quantify these parameters based on their importance to spatter generation.
@@
Figure 14: Spatter prediction on FLOW-3D simulation generated from testing FLOW-3D dataset at 400 W 2 m/s on ML model that was trained using OpenFOAM dataset. Region masked with black are identified as location predicted as spatter (a) prediction at timestep 65 $\mu s$ (b) prediction at timestep 165 $\mu s$ (c) prediction at timestep 265 $\mu s$.
3.3 Spatter and Melt Pool Classification via ML Model.
To investigate the performance of the OpenFOAM data sets on ML models, features included the x-axis, y-axis, z-axis, temperature, velocity, x-velocity, y-velocity, z-velocity, density, and pressure are used as the model input. Consistently high accuracy was observed across all ML models, with training accuracy ranging from 98% to 100 % and testing accuracy ranging from 97% to 99%. The high accuracy was maintained across all calculated score metrics (Accuracy, F1, Balanced Accuracy, ROC_AUC) as shown in Figure 10(a-d) The lowest test accuracy of 97% was observed with the XGB and SVC models. The closeness of the training and testing accuracy indicates that the ML models are neither under- or over-fitting. The importance of the characteristics used in the ML test, as shown in 10e, followed the trend: z-axis $>$ pressure $>$ z-velocity $>$ y-axis $>$ temperature $>$ velocity $>$ x-velocity $>$ y-velocity $>$ density. The highest contribution of the z-axis to accuracy is attributed to the wide separation observed between the spatter and the melt pool when the z-axis is paired together with itself, or other parameters, as shown on the histogram, KDE, and scatter plots as shown in 7.
@@
The methodologies of our data extraction and combination of parameters for this classification task is promising. These techniques can serve as a benchmark in predicting spatter or developing reduced order model package for spatter generation either with desired simulation packages or collected experimental measurements. Modeling LPBF in FLOW-3D generates a dataset that is well-suited for analyzing the melt pool rather than the spatter. Some parts of the melt pool dataset generated by FLOW-3D might have properties similar to those of spatter or meet the threshold required for spatter ejection. However, these locations remain part of the melt pool in the simulation because the physics module in FLOW-3D uses simplified assumptions for mass transfer between different phases, lacking the capability to accurately simulate spatter. Adapting the current ML model that was trained on OpenFOAM dataset will necessitate removal of spatial components (x, y, z) from the features that will be used to train and test ML models. On training and testing OpenFOAM dataset without spatial components, high accuracies were still achieved with slight decrease by  2% across the model performances in comparison to when spatial components were present as shown Figure 11. Significant decrease of 8 % was observe with the KNN model. The accuracy varies from 97 % to 95 % which are within the same range. High accuracy of the ML model with the features without spatial component qualify any one of the model for predicting spatter on FLOW-3D package. To ensure the models are not overfitting, we compared both the training and testing score as shown in the Figure 11. The largest difference between train and testing accuracy that was observed is 0.1 with respect to the KNN model, while the rest of the models are less than 0.05. These slight changes are small to consider the investigated model as not overfitting. The trend observed in the feature importance studies without spatial components, as shown in Figure 11e follows the trend of pressure $>$ z-velocity $>$ temperature $>$ velocity $>$ x-velocity $>$ y-velocity $>$ density. Evidently, pressure is not only the most significant parameter driving the dynamics in the melt pool, but also dictates possibility of spatter ejection as its magnitude is confined within a narrow range as shown in Figure 5. Strong importance observed with respect to z-velocity can be understood from the tendency of a body that possesses high z-velocity to travel upward, and we observed a narrow range for melt pool z-velocity, which is a subset of spatter’s z-velocity. Also, observing temperature having less importance than the z-velocity despite its direct relationship with pressure according to Clausius–Clapeyron relation could be attributed to the saturation that is reached such that increment in the temperature will result in boiling of the material which will not lead to spatter ejection, but more plume generation. In addition, Khairallah et al. [68] demonstrated the critical role of recoil pressure and Marangoni convection in shaping melt pool flow and influencing the formation of defects such as spattering, denudation, and pores. They found that as temperature rises, the recoil pressure increases, overcoming surface tension and generating depressions and spatter. High contribution of z-velocity to the spatter ejection could also be understood from the parabolic shape of the melt pool which is formed as the laser traverse in x-axis direction. Hot fluid will build up in the direction of the laser, which can break up to create spatter due to its elongation as shown in Figure 6a, via shear thinning. Additionally, spatter breakage can further intensify the backward dispersion of material and energy that are in form of a wave which can travel through the high gradient region of the melt pool, creating z-velocity component that could enhance ejection. The backward dispersion of material and energy in the melt pool can be attributed to the combination of other factors such as pressure, temperature gradient, and the vibration in the melt pool which can also influence other properties like viscosity and surface tension to cause Marangoni effect and promote melt pool flow, elongation and ultimately fluid ejections. To gain further insight to the distribution of ML model predictions, we constructed a precision recall table as shown on Figure 11f and 10f for predictions with spatial components included in the features and with no spatial components respectively. We observed our model having the capability in predicting most of the points with the melt pool and the spatter. The wrong prediction observations for the model with the presence of spatial component features and non-spatial component features is associated with regions that were melt pool but were falsely predicted as spatter and three and six respectively in number.
3.4 Explainable AI via SHAP Analysis.
The shap.TreeExplainer was used to calculate SHAP values for the test dataset, focusing on the contribution of features to the prediction of the positive class (spatter). Comparison of SHAP importance and ML importance ensures consistency in feature relevance across different methods. A summary bar plot was generated to visualize the mean absolute SHAP values, which indicate the average magnitude of feature impacts on the model’s output. The SHAP summary bar plot offers a high-level overview of the importance of the features. Figure 12a and Figure 13a present the feature important from the SHAP analysis for the experiment with and without spatial components respectively, in the datasets. The x-axis represents the mean absolute SHAP values for each feature, indicating the average magnitude of a feature’s impact on the model’s output. Features are ranked in descending order, making it easy to identify the most influential features driving predictions. The SHAP feature importance rankings are consistent with the feature importance rankings derived directly from the ML model in Figure 10e and Figure 11e indicating consistency between the model’s internal behavior and the SHAP explanation.
The SHAP summary plot with a color bar adds depth by visualizing the distribution of SHAP values for individual samples. Each point corresponds to a sample, with its position along the x-axis representing its SHAP value and its color encoding the feature value (e.g., high or low). This plot reveals both the importance and directionality of features, indicating patterns and interactions in the data. The farther the SHAP values deviate from zero, the stronger the influence of the feature on the model output, which was observed to follow the trends observed in 10b and Figure 11b. Furthermore, high values (red) are predominantly associated with positive SHAP values in high-influencing features such as Z-Axis, pressure, Z-Velocity, velocity magnitude, indicating their ability to increase the likelihood of spatter predictions. In reverse, a high value (red) of variable such as temperature and Y-axis tends to negative contributions to spatter generation.
To further understand the effect of changing the value of a particular feature on the prediction of the model, PDP plots were generated for each feature as shown on the Figure 12(c-l) and Figure 13(c-i). The y-axis of the PDP shows the average predicted probability of the positive class (spatter). In Figure 12c, we observed that the Z-Axis value has a significant threshold effect from -10 $\mu m$ to 40 $\mu m$ suggesting this range is critical for the model’s predictions, while beyond  100 $\mu m$, the impact saturates. The extreme negative and positive values of the Y-axis strongly influence the model output, while values close to 0 have a minimal effect that indicates the symmetric behavior of the Y-axis to influence the predictions of the model with respect to the presure (Figure 12d) and (Figure 13c), below  0.2 MPa, the partial dependence remains flat, indicating negligible impact on predictions. A sharp increase in partial dependence is observed as the pressure increases from approximately 0.1 to 0.2 MPa, where a significant positive impact is observed on the model predictions. The partial dependence is nearly flat and stable around 0.35, indicating that pressure in this range has minimal influence on the model’s predictions. The curve plateaus around 0.55 to 0.60, suggesting that increasing pressure beyond this point does not have a substantial additional impact on the predictions. A significant increase in Z velocity was observed from 2 m/s to 5 m/s (Figure 12f), while the velocity magnitude initially increased at 2.5 m/s, then almost stabilized to 10 m/s where it continued to increase to 17 m/s (Figure 12 g). Close partner was observed in the PDP plot of features with the ML experiment, where there are no spatial components in the input features in Figure 13. Steady reduction in temperature was observed to favor spatter generation (Figure 12h and Figure 13e) which could be attributed to spatter generation occurring mostly away from the location of the laser, as the fluid swirl backward due to the Marangoni recirculation effect. Along the X-Axis (the laser traveling direction) in Figure 12j, SHAP value decline at the initial stage then remain almost steady at the higher value X-Axis. Although the SHAP values of features like Y-Velocity, density, and X-Axis are small, it is likely that they promote the accuracy predictions in our studies due to their tendency to interact with features with high SHAP values.
3.5 Spatter Prediction on a Commercial Package ”FLOW-3D”.
@@
Spatter prediction on FLOW-3D is promoted by two factors: (1) the close alignment between the characteristics of the melting pool dimensions and the temperature distribution of OpenFOAM and FLOW-3D and (2) the high precision of the ML model in classifying the spatter from the melting pool. Upon testing our model on FLOW-3D, voxels that match the spatter property were identified and masked with black points as shown in Figure 14. At the early timesteps of the FLOW-3D simulation, spatter predictions were predominantly observed at the front of the melt pool (Figure 14a), then at the last time steps (Figure 14b and Figure 14c), the spatter tends to occupy the backward side of the melt pool, this is similar to the observed OpenFOAM simulation for spatter generation as shown in Figure 9 and the experimental result reported by Yin et al. [5]. Note that conditions for spatter generation were not met at the location of the hotspot of the melt pool because of the high temperature that could be causing fluid evaporation. Consequently, the high accuracy outcome of the Random Forest (RF) model underscores the efficacy of our developed methodology for the reduced order task. This methodology can enhance the robustness of commercial packages by synergizing them with machine learning (ML) models to account for complex and computationally expensive phenomena, such as spatter modeling. An alternative application of the developed methodology is testing datasets of different process parameters generated from commercial software, such as FLOW-3D, towards the development of a process map.
@@ 3.6 Spatter Process Map
Figure 15: Process map generated via hybrid of OpenFOAM trained ML model on FLOW-3D dataset to understand relationship between spatter formation and process conditions. The process map was created using 283 FLOW-3D simulations with different combinations of power and velocity.
Display of spatter quantification with varying power/velocity combinations will provide access to a suitable operating window with minimal defects for part production. Running OpenFOAM for additive manufacturing is computationally expensive due to a lack of simplifying assumptions in the physics it is configured to solve as compared to FLOW-3D, that has a devoted module for AM. On the same machine, both FLOW-3D and OpenFOAM require run times of 4 and 72 hours respectively to generate LPBF solution on 5 $\mu m$ mesh size using Intel(R)Xeon(R) Gold 6230R CPU @ 2.10GHz, 2095Mhz, 26 cores. Building a hybrid of OpenFOAM and FLOW-3D via the high accuracy ML model as a reduced order model is attractive for large dataset generation for the development of spatter process map. Figure 15 is generated by testing 281 dataset consisting varying power/velocity combinations of FLOW-3D experiments on OpenFOAM trained ML model. The red and green lines overlay on the plot are extracted from a previous study on porosity by Rollett et al. [12] indicating boundary within which keyhole and lack of fusion (LOF) respectively were prominent. In-between the boundary is the window identify as process window where keyhole and LOF are minimal. These lines are overlaid on the spatter count process map to relate the generated spatter count at a specific power/velocity combination to the nature of porosity that is likely to be present with the aim of identifying region on the process map where minimal effect of porosity and spatter can be seen during part building. As shown in Figure 15, the spatter quantification follows a predictable trend within processing parameter space for LPBF, except in a region between 500 - 700 $mms^{-1}$ where spatter count is consistently low across the power range investigated. This is also clearly demonstrated in Figures A.11 and A.12 of the Appendix. It was observed that higher spatter was generated at high power and low velocity, which is the region associated with keyhole formation. Conversely, at low power and high speed, less spatter was observed, corresponding to the region characterized by lack of fusion [12]. Based on the process conditions, we observed the trend followed by the volume of spatter generation as high power/low speed $>$ high power/high speed $>$ low power/low speed $>$ low power/high speed. The observed trend can be attributed to the intensity of the interaction between the material and the laser, leading to variations in melt pool size, and the magnitude of mass and energy transport. Consequently, high power and low velocity produce the largest melt pool with the most intense transport phenomena and kinetic effects, resulting in the highest spatter ejection. In contrast, low power and high velocity create the smallest melt pool with the least intense effects, resulting in minimal spatter ejection. Certain locations within the safe operating window for porosity (situated between keyhole formation and lack of fusion) [12] intersect with areas of low spatter count, as shown in Figure 9. These intersections make those regions promising for part printing. The reason for the observed low spatter volume at a velocity between 500 - 700 $mm/s$ throughout the investigated power range is probably the transition between a kinetic regime and a diffusion-limited regime resulting from a change in the morphology of the melt pool. This transition creates a balanced environment in which the dynamics of the melt pool is stabilized, leading to reduced ejection of spatters due to moderate vaporization, stable formation of the melt pool, and efficient heat distribution.
The analysis of the study by Zheng et al. [69] further supports the transition observed on the process map from 400 - 700 mm/s. They examined the formation and evolution of the plume and its impact on the generation of spatter in stainless steel at varying scan speeds. ( 200 mm/s, 400 mm/s, 500 mm/s, 1000 mm/s, and 2000 mm/s). They believe that spatter generation is more likely to depend on the stability of the plume throughout the experiments. They discovered that at lower scan speeds (200 mm/s), the vapor plume was highly unstable because of the high energy density, with recoil pressure strong enough to eject droplets from the melt pool. As the laser scan speed increased to 400 – 500 mm/s, spattering and plume formation barely occurred. Further increasing the speed to 1000 mm/s and 2000 mm/s caused the melt track to become unstable, showing periodic swelling connected by narrow bottlenecks of ripples along the melt track, a result of the Plateau-Rayleigh capillary instability [70]. When the scan speed exceeds a certain threshold (500 mm/s in our study), the vapor plume tilts backward relative to the scan direction. This tilt is believed to result from a shift in the laser-material interaction interface from the bottom to the front wall of the depression or keyhole. The resulting backward vapor flux is thought to further influence the fluid dynamics within the melt pool, increasing the susceptibility of the melt track’s end to open-pore formation.